Outline for E/D Nights UT February - March 2021

Nights shared with PETS

  • DONE UT Feb 3 (local Feb 2) - handover from PEPSI/PETS at 23:45 - until morning twilight (lost to weather)
  • DONE UT Feb 4 (local Feb 3) - PEPSI/PETS 21:00-03:30 (local time) (lost to weather)
  • DONE UT Feb 8 (local Feb 7) - 100% PEPSI/PETS
  • DONE UT Feb 9 (local Feb 8) - PEPSI/PETS 20:30-03:00 (local time)
  • DONE UT Feb 14 (local Feb 13) - PEPSI/PETS 21:30-04:15 (local time)
  • DONE UT Mar 3 (local Mar 2) - PEPSI/PETS (twilight) - 23:30 UPDATED-UPDATED
  • DONE UT Mar 4 (local Mar 3) - PEPSI/PETS 20:40-02:01 (start of PETS is 22:00, handover 30min before to set up) UPDATED-UPDATED
  • UT Mar 14 (local Mar 13) - PEPSI/PETS 20:15-02:16 (start of PETS is 21:45, handover 30min before to set up) UPDATED-UPDATED
  • PETS has priority (backup with MODS or LBCs suggested) - do not use LUCI or LBTI before PETS time during the night
    • Parts for SX M3 repair due mid-March

Nights shared with LBTI or Regular PEPSI

Nominally a 50-50 split
  • DONE Share with PEPSI - UT Feb 25 (local Feb 24) - PEPSI first half - night center 00:25
  • DONE Share with LBTI - UT Feb 26-Mar 2 (local Feb 25-Mar 1)
  • DONE Share with LBTI - UT Mar 5 (local Mar 4)

LBTI Requests

  • UT Feb 26 (local Feb. 25): First three hours of the night plus second half.
  • UT Feb 27 (local Feb. 26): First two hours of the night plus second half.
  • UT Feb 28 (local Feb. 27): First half or first few hours of the night plus second half if part of the D time the previous nights is lost.
  • UT Mar 1 and Mar 2 (local Feb. 28 and Mar. 1): First half of the night.
  • UT Mar 5 (local Mar. 4): First half of the night.

Engineering Tasks

  • AO On-sky checkout
    • Bin 1 and 3 Full AO - 2hr
      • multiple PAs
      • multiple elevations
    • Additional tests - (E. Pinna, D. Miller, B. Rothberg)
      • closed loop on AO ref stars (on-axis) to collect SR and telmetry
        • one or more stars in the following categories
          • DONE (2 completed) 14.5 < Rmag < 15.5 Fast Bin 3
          • DONE (2 completed)15.8 < Rmag < 16.5 Slow Bin 3
          • DONE (1 completed) 16.51 < Rmag < 17.0 Fast Bin 4
          • DONE (2 completed) 17.0 < Rmag < 18.0 Slow Bin 4
          • Rmag > 18.0 (as far as we can go....) - only went down to WFSMag=17.9
    • Bin 2 ESM - multiple PAs - 1 hr each - any conditions
      • M67 H-band AO Ref R=12.2 - PA 0,60,120,180,240,300
        • Sets at 4:15am
      • M92 H-band AO Ref R=9.4 - PA 0,60,120,180,240,300
        • Visible 3:30am through twilight
    • Hoag's Object - On-axis AO Ref star (using core of the galaxy) WFSmag estimate = 14.7 - 1hour shutter time (2nd half object)
  • DONE MODS1 Active Optics parameter change (JMH) (done for MODS1 20210209 UT, done for MODS2 20210210 UT)
    • needs ~1 hour on-sky with MODS1 collimating with ~3 different guide stars (don't care what MODS is looking at)
    • priority=medium-low, seeing=average, clouds=don't care
  • DONE LUCI Active Optics parameter change (JMH) (done 20210224 UT)
    • needs ~1 hour on-sky with LUCIs collimating with ~3 different guide stars (don't care what LUCI is looking at)
    • priority=low, seeing=average, clouds=don't care
  • DONE LBC Pointing Hysteresis and Settling Test (JMH/OPK)
    • needs 1 hour max in any conditions where we are open
    • priority=medium-low (interesting for TMS)
    • see detailed plan at the bottom of this page
  • DONE LBC bino observations with active TMS control
    • 3 hours observing some target rising or setting
    • priority=medium (Not ready before mid-Feb)
    • schedule for Feb 24-Mar 4 D/E blocks
  • DONE MODS1+LBCR mixed-mode on-sky test
    • 0.5-1 hr
    • Run a long MODS observation with MOS-type alignment (i.e., dtheta, dx, dy) and a simultaneous LBCR dithered sequence.
      • Make sure we can stop LBCR, refocus/recollimate, and then load & play a new LBCR OB.
    • priority=high (for an LBTB program), seeing=poor, clouds=don't care
  • MODS Field Correction check for IT 8315
    • 1-1.5 hours in seeing better than 1.4 arcsec, clouds don't matter much.
    • priority=medium (problem may have been with us for a year)
    • run IDL task field_collect (same as measuring field aberrations: ComTools#Field_Data_Collection)
      • do LDG and RDG in parallel (two IDL windows)
      • Use 6-8 stars (some fields only have 6 stars; start the list a second time, if you are under an hour)
      • Stone_G is well placed at RA=10:17, and other Stone / BS fields are available if earlier or later.
  • Pointing Model Data Collection
  • Standard Stars (only in photometric conditions)
    • MODS primary spectrophotometric standard in all modes
      • The table of MODS primarty spectrophotometric standards is here
      • Scripts for acq and observation (all modes) are in ~modseng/modsScripts/modsSpecPhot
    • MODS imaging standards (this is better done later - the fields are only just rising at the end of the night in March)
    • LBC imaging standards
    • LUCI Standards
      • Scripts: /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/LUCI_Phot_ZP
      • Best to pick one star and sample it over different airmasses
      • Otherwise, make sure to sample airmasses over delta-airmass > 0.33

LBC Pointing Hysteresis and Settling Test Procedure

  • Preparation:
    • Prepare targetless OB to take short 1-5 sec subframed exposures with minimum readout time. Set Nexp to 999 to make it continuous.
    • Set LBC to NOT "Use Telescope", and NOT "Use Mirror"
    • Deploy LBCs and authorize. LBCR required, LBCB optional
  • Execution:
    • Use IRAF/IDL to send TRACK preset to a field near the pole at EL=50 (DEC=+72)
    • Run dohybrid for focus and collimation (no need for copointing)
    • Play the OB for 2 exposures just to see that it works. Then cancel.
    • Start TMS taking continuous data (optional).
    • Move telescope to zenith, and sit there at least 3 minutes.
    • Play the LBC OB to start taking continuous short exposures.
    • Use IRAF/IDL to send TRACK preset to a field near the pole at EL=50 (DEC=+72)
    • Sit there for 10 minutes continuing to take exposures.
    • Slew to hold at EL=30.1, and sit there 3 minutes.
    • Use IRAF/IDL to resend TRACK preset to a field near the pole at EL=50 (DEC=+72)
    • Sit there for 10 minutes continuing to take exposures.
    • Stop the LBC OB.
  • Afterward:
    • Record from the event log, the times of preset completion.

Science Requests

  • 2020 SO - 2 hours (1 hour each night) - No longer visible after UT Feb 4
  • Solar System objects recovery - 2-3hrs - LBCs - (Veillet)
  • DONE Spectroscopic followup of IMBH in low-metallicity dwarf galaxy - MODS - 1.5 hours (Rothberg)
    • RA 16hr Dec +33 - target is visible last 3 hrs of the night.
    • Required Conditions: Seeing < 1.2", thin clouds acceptable
    • Darker skies better
  • mid-z ULIRGs MODS - 2hours per target - 2hours allocated - may request additional 2 hours for 2 target total (Rothberg)
    • FF0234 - highest priority
    • FF0238 - 4th priority
    • FF0937 - 3rd priority -if FF0234 or FF1106 cannot be done - this is OK for any Moon (2nd prority in bright time)
    • FF1106 - 2nd priority if FF0234 cannot be done - requires darker skies
    • Seeing < 1.2", thin clouds OK, darker skies better
  • DONE V1112 Per - Nova - MODS - 1hr w/overheads
    • 04:29:18.84 +43:54:23.2
    • V=16, R=14.5
    • LD-2021A-0005
    • Any Conditions
  • DONE ALES HD33632 brown dwarf companion observation
    • 05 13 17 +37 20 14
    • Directors time from December, lost to weather and M3 issue
    • 3h at the beginning of the night
    • Feb 25 or Feb 26, later we won't get enough field rotation
    • Weather: clear, <1.2-ish arcsec seeing
  • DONE LMIRCam M band HD33632 brown dwarf companion observation
    • 05 13 17 +37 20 14
    • Directors time from December, lost to weather and M3 issue
    • 2h at the beginning of the night, not the same night as ALES observation
    • Febb 25 or 26, later we won't get enough field rotation
    • Weather: clear, <1.5-ish arcsec seeing

Poor Weather Backups

  • ESM Data collection
    • Bin 2 and Bin 4 - M67 (sets at 4:15am) or M92 (rises at 3:30am)
    • Run through JHK, with bright and faint AO Ref stars
  • ICRF3 backup programs - LUCI-Bino - Rothberg - 0.5hr per objects - any conditions
  • Velocity Dispersion of local mergers - LD-2019B-005 - 1hr per target bino-MODS
    • DONE U4079
    • DONE N3310
    • DONE N2744
    • N6598 - all up 2nd half of night
    • any conditions - go longer if poorer conditions
  • Velocity Dispersions of local mergers - LD-2019B-005 - 1hr per target bino-LUCI
    • DONE NGC 3310 10 38 45 53 30 05 CWL=2.27 GS R=12.5
    • Arp 156 10 42 38 77 29 41 CWL=2.30 GS R=14.1 (can only observe between 22hr-02hr to get telluric, if telluric obtained before 02hr then can continue past)
    • NGC 4004 11 58 05 27 52 44 CWL=2.27 GS R=14.2
    • UGC 9829 15 23 01 -01 20 50 CWL=2.30 GS R=13.0
    • Any conditions - go longer if poorer conditions
  • MODS observations of RLQ/RQQ QSOs
    • script location: /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/Rothberg_Science/QSOs_MODS
    • Any conditions, list is not prioritized
    • LD-2019B-004
      • He1029-1401
      • PG 1307+085
      • PG 1416-129
      • HB89 1635+119
  • QSO Test Targets for H-band Library Program - LUCI-Bino
    • LD-2019B-003
    • ~45 min per target with overheads
    • Script Location: /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/Rothberg_Science/H-band_Library/
    • DONE PG 1444+407
    • PKS 1302-102

Script Locations

  • NEW - scripts for 2021A are located here:
    • /home/lbto/Service_Observing/2021A_D_E/
    • scripts for this page are in sub-directory:
      • Feb_Mar_2021

M3 Amplifier Fault Workaround

  • If the OSSGUI reports an amplifier fault they should power cycle the UMAC (power off, count to 10, and power on) and then attempt to move the unit.


General Night Plans

UT Feb 3

  • Reconfigure for LUCI-Bino, LUCI-1 SX authorized
  • OSCO - 0.5 hr
  • Bin 2 and Bin 3 tests Full AO, multiple PAs and elevation - 1.5 hr
  • 2020 SO LUCI-1 AO - 1.5 hr
    • Bin 3 working required for proceeding to 2020SO - otherwise force to Bin 4
  • ESM tests Bin 2 - multiple PAs and elevation - 1hr
  • ESM data collection - 1.5 hr
  • If SX M3 fails to move into position - reconfigure for MODS
    • IMBH candidate (Rothberg) MODS - 1.5hr
No Chance to Open - no tasks done

UT Feb 4

  • Configure for bino-MODS at sunset - unless clear skies - if clear - take LBC skyflats
  • Reconfigure for MODS if LBC skyflats taken 0.3hr
  • LBC-R/MODS-1 Mixed mode tests - 1hr
  • MODS1 Active Optics parameter change (JMH) - 1hr
  • Reconfigure for PEPSI - PETS
  • Reconfigure to LBCs - 03:30hr AM
    • 2020 SO (CV)
No Chance to Open - no tasks done

UT Feb 9

  • Configure for bino-MODS at sunset - unless clear skies - if clear - take LBC skyflats
  • Reconfigure for MODS if LBC skyflats taken 0.3hr
  • LBC-R/MODS-1 Mixed mode tests - 1hr
  • MODS1 Active Optics parameter change (JMH) - 1hr
  • Reconfigure for PEPSI - PETS
  • Reconfigure to LBCs
    • Hysterisis Tests
    • TMS Testing
On-sky - Mixed-Mode tests executed, Active Optics tests executed, Hysterisis tests executed

UT Feb 14

  • Twilight to 21:30 local time - LBC or MODS
  • Reconfigure from PEPSI t0 MODS at 04:15
    • IMBH Target - 1.5 hours (Rothberg) -MODS
No Chance to Open - no tasks done

UT Feb 25 (local Feb 24)

  • Night split 50/50 w/PEPSI science - D/E-time 2nd half - Night Center 00:33 - handover 00:28
    Moon is up effectively all night (94.5%)
  • AO Engineering Time - highest priority - (LBTO AO and Arcetri)
    • LUCI-SOUL AO tasks - test 6 different PAs per object
    • LUCI-SOUL Arcetri tests
    • LUCI-SOUL ESM tasks - 1hr - only if no other full AO
      • Hoag's Object - on-axis AO Ref star (galaxy core as AO Ref star R=13.44 predicted WFSmag =14.7)
  • if cannot switch to LUCI
    • LBCs - TMS testing if AO completes early
    • LBCs or MODS for science
    • LBCs or MODS Eng tests
Open for 2nd half - variable seeing, thick cloud cover at times

UT Feb 26 (local Feb 25)

  • Night split with LBTI - night center 00:32 - handover 00:27
  • Moon is up effectively all night (98.4%)
  • No LUCI-AO support needed
  • 1st half D/E-time
    • Twilight (19:15hr)-22:15 - HD 33632 - LBTI w/ALES
    • MODS Field Correction check for IT 8315 - 22:30-00:27 - MODS
      MODS Science
      • V1112 -1hr
      • NGC 3310 -min 0.5hr
  • 2nd half LBTI
Poor seeing 1st half - no LBTI data, backup MODS executed

UT Feb 27 (local Feb 26)

  • Night split with LBTI - night center 00:32 - handover 00:27
  • Moon is up all night (99.8%)
  • Roll-over previous night plan if lost to weather
  • No LUCI-AO support needed
  • 1st half D/E-time
    • Twilight (19:15hr) - 21:15 - HD 33632 - LBTI w/LMIRCAM
    • LBCs - TMS testing and/or Solar System recovery (CV) - 21:35-00:27
  • 2nd half LBTI
LMIRCam Data obtained,, TMS testing executed (poor seeing)

UT Feb 28 (local Feb 27)

  • Night split with LBTI
  • Moon is up all night (98.6%)
  • Roll-over previous night plan if lost to weather (or night before that)
  • Plan TBD based on losses or completion of programs prior to this night
  • No LUCI-AO support needed
  • 1st half D/E time:
    • Twilight (19:15hr)-22:15 - HD 33632 - LBTI w/ALES
    • Science: V1112 -1hr (original data taken in 2-3" seeing, pending PI check)
    • MODS Field Correction check for IT 8315 - 22:30-00:27 - MODS
    • Science: UGC 4079 (to fill time until handover if needed) - MODS
  • 2nd half LBTI
ALES obsrevations attempted, lots of AO-LBTI issues, closed for winds at switch to MODS - never reopened

UT Mar 1 (local Feb 28)

  • Night split with LBTI - night center 00:32 - handover 00:27
  • Moon rises 19:57 - through twilight (94.6%)
  • LBTI 1st half
  • D/E 2nd half
    • LUCI-SOUL AO On-sky Testing - (LBTO AO and Arcetri)
No LUCI-AO-SOUL - M3 SX would not move into position - high winds prevented any data collection after handover

UT Mar 2 (local Mar 1)

  • Night split with LBTI- night center 00:32 - handover 00:27
  • Moon rises 21:05 - through twilight (88%)
  • LBTI 1st half
  • D/E 2nd half
    • LUCI-SOUL AO On-sky Testing
LBTI aborted due to UBC vacuum event - MODS backup science (poor seeing) done - MODS2 offline - LUCI backup science done 2nd half - 1 hr of AO attempted - poor/variable seeing

UT Mar 3 (local Mar 2)

  • Night split with PETS
  • D/E-time: 23:40 (includes reconfig time from PEPSI)-twilight - Moon rises 23:22 - through twilght (79%)
  • LUCI-SOUL AO On-skytesting 2nd half
LUCI-1 SOUL-AO testing went well, collected data as planned. The only significant issue is the worsening capabilities of the DIMM. It is becoming more difficult to lock on to stars. Anything fainter than about V~2mag seems to cause problems.

Arcetri testing depends heavily on data from the DIMM, this resulted in some hunting around and moving around to find decent fields within 20 degrees of 2nd mag or brighter star (it even had trouble with Arcturus from time to time!)

UT Mar 4 (local Mar 3)

  • Night split with PETS
  • D/E-time: twilight-20:40 (30min early to reconfig and setup PEPSI) MODS or LBC - Moon is down
    • mid-z ULIRG FF0937 - ~ 1.5 hour w/overheads
  • Handover to LBTO at 02:26
  • D/E-time: 02:26-twilight (includes reconfig time from PEPSI) - Moon is up (69%)
    • Rothberg IMBH candidate - MODS-1 only is OK - will collect more data following night to make up for loss of MODS-2
Opened brielfy for PEPSI/PETS - all D/E time lost due to winds, humidity, and eventually snow

UT Mar 5 (local Mar 4)

  • Night spliit with LBTI - night center 00:31 - handover 00:36
  • Moon rises 00:31 - through twilight (58%)
  • LBTI - 1st half of night
  • D/E-time 2nd half Science Programs
    • Rothberg mid-z ULIRG FF0937 - 00:36-02:36
    • Rothberg IMBH candidate 02:40-04:10

UT Mar 14 (local Mar 13)

  • No Moon up
  • D/E-time twilight-21:15
  • PEPSI/PETS 21:45-02:55
  • D/E-time 02:55-twilight
    • Possible ESM use last few hours of night

UT Mar 29 (local Mar 28)

  • Full D/E Night


-- BarryRothberg - 20 Jan 2021
Topic revision: r76 - 05 Mar 2021, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback